# Quantifying the Uniqueness of Donald Trump in Presidential Discourse

### v1

@@ Abstract
Does Donald Trump speak differently from other presidents? If so, in what ways? Are these differences confined to any single medium of communication? To investigate these questions, this paper introduces a novel metric of uniqueness based on large language models, develops a new lexicon for divisive speech, and presents a framework for comparing the lexical features of political opponents. Applying these tools to a variety of corpora of presidential speeches, we find considerable evidence that Trump’s speech patterns diverge from those of all major party nominees for the presidency in recent history. Some notable findings include Trump’s employment of particularly divisive and antagonistic language targeting of his political opponents and his patterns of repetition for emphasis. Furthermore, Trump is significantly more distinctive than his fellow Republicans, whose uniqueness values are comparably closer to those of the Democrats. These differences hold across a variety of measurement strategies, arise on both the campaign trail and in official presidential addresses, and do not appear to be an artifact of secular time trends.
@@ 1 Introduction
President Donald Trump’s distinctive speech patterns have attracted a considerable amount of media attention.111“The version of English he speaks,” notes Kurt Anderson, “amounts to its own patois, with a special vocabulary and syntax and psychological substrate” [1]. Meanwhile, Andrew Romano makes the point a good deal more plainly: “Donald Trump doesn’t talk like other presidents” [2]. And with good reason. More than just rhetorical flourishes, patterns of presidential speech are constitutive of what Miroff calls the “spectacle” of presidential leadership [3]. Moreover, as both a populist and insurgent within the Republican Party, Trump’s chosen ways of talking are specifically intended to set him apart from other U.S. politicians and enhance his political appeal [4].
A growing body of scholarship examines the characteristics of Trump’s speech patterns, both qualitatively and quantitatively. According to Hart [4], Trump uses language that is explicitly intended to evoke an emotional response from his audience. And other scholars argue that Trump’s speech patterns reveal his autocratic ambitions [5], anti-democratic views [6], a willingness to condone political violence [7], a commitment to populism [8], and low levels of analytic thinking [9]. Quantitative approaches used in these studies are primarily based on lexicons (e.g., [9]) and sentiment classification (e.g., [10]), which necessarily ignore inherent contextual information in speeches. As a result, it is difficult to capture patterns suggested by qualitative studies — e.g., Jamieson and Taussig’s observation that Trump distinguishes himself by being “spontaneous and unpredictable” [6].
@@
Leveraging recent advances in large language models [11], we develop a new quantifier of uniqueness by directly measuring how unpredictable a speech is given the preceding context. In addition, we zoom in on a prevalent yet understudied construct, divisiveness [8]. We operationalize divisiveness as language that is intended to impugn and delegitimize the speaker’s target, and we develop a new lexicon for such speech. Furthermore, we leverage mentions of opponents, especially prevalent in presidential debates, and introduce a comparative framework for the portrayal of political opponents involving lexical features.
We use our proposed tools to analyze a large and diverse corpus of presidential speech. In doing so, we are able to both characterize the overall distinctiveness of Trump’s speech patterns among recent presidential candidates and examine specific qualities that previous research has overlooked. On the campaign trail and in official presidential addresses, we show that Trump is unlike any other recent president. Like previous scholars [12, 13], we show that Trump tends to communicate in shorter, more simplistic sentences. Unlike previous research [9], we find that Trump speaks in ways that are different from all modern presidents, including his immediate predecessors and successor.
Some notable findings from our analysis include his use of distinctive descriptors for his political opponents. Indeed, Trump’s use of divisive and antagonistic language and his tendency to target political opponents distinguish him from all other presidential candidates. Furthermore, Trump stands apart from all his fellow Republicans, whose uniqueness scores are comparatively closer to those of the Democrats. We also document Trump’s patterns of repetition for emphasis.
@@ 2.1 Data
Our research investigates three sources of political speech: presidential debates in general elections since 1960, State of the Union (SOTU) speeches since 1961, and a sample of campaign speeches assembled by the American Presidency Project [14]. Debates and SOTU speeches are generally standard across presidents and time, so all available documents are included in our main analyses. [^2] Additional data were used for model training. For details, see §A.1 and §A.3 for details. Publicly available campaign documents, however, are imbalanced and not comprehensive over the same time frame, so these corpora are limited to speeches delivered within one month of Election Day since 2008. Table 1 provides statistics of the final datasets for which we present results. Of these datasets, Donald Trump spoke in five debates (3,610 sentences), four SOTU addresses (1,471 sentences), and 28 campaign speeches (7,488 sentences) between the years of 2016-2020.
Fig. 1 shows the distribution of sentence lengths from each speaker, across the datasets. Further details of how we collected these datasets can be found in §A.1. From the outset, however, we note that Trump tends to speak in markedly shorter sentences than other presidents do (except Biden). Whereas Trump’s sentences range from 10.4 to 14.5 words in the three data sources, the overall averages ran from 17.6 to 24.4 words. And when comparing presidents within each data source, Trump registered the single lowest number of average words per sentence among all presidents.
Figure 1: Distribution of sentence lengths among all speakers, across debates, SOTU, and campaign data (error bars represent 95%-confidence intervals). Donald Trump tends to use the shortest sentences on average.
@@ 2.2 Quantifying the uniqueness of political speech
We quantify the uniqueness of political speeches using three complementary approaches: (1) a novel metric based on large language models (LLMs); (2) a new resource of lexicons for divisive speech; and (3) a comparative framework for portrayal of political opponents involving lexical features. Together, these three approaches allow for a robust and multifaceted comparison of speech patterns and shed insights on Trump’s speech patterns both beyond lexical uses and from the perspective of divisiveness. We briefly introduce the intuition behind each approach below.
LLM-based uniqueness
@@
Large language models, like the GPT family of models, have received widespread attention for their abilities to statistically characterize the complex structures of natural language text. LLMs can measure the predictability of text by calculating the likelihood of the next word or sequence of words in a given context, to produce a measure known as perplexity. Perplexity is typically used to evaluate the quality of LLMs [15, 16]. Bits-per-character (BPC) has been used in lieu of perplexity to control for length [17].
@@
Contributing to this literature, we propose a metric of “uniqueness” based on the ability of LLMs to estimate the probabilities of word sequences, and we then use these estimates to compare political speech from various presidential candidates. Specifically, from a pool of candidates, we determine how likely the speech of one speaker is to be produced by the others. A positive value for one candidate’s sentence uniqueness, therefore, suggests that other candidates are unlikely to say it. The advantage of this metric is its consideration of the preceding context of given speech. The precise meaning of the scores, however, may not be interpretable, especially since they only allow for comparisons within a specified pool of speakers. Technical details on the construction of this metric are included in §A.3 and §A.4.
Divisive speech lexicon
@@
To analyze the actual content of language used by candidates, a “divisiveness” lexicon is created and applied to each dataset. We define language as “divisive” if it intends to impugn and delegitimize the speaker’s target, e.g., by attacking their intelligence, integrity, or intentions. Examples of divisive accusations include “racist”, “dishonest”, “corrupt”, or “ridiculous”. Such labels are expressly designed to put the target on defense and accentuate differences and distance between parties. Our definition of divisive is distinct from other commonly analyzed constructs such as political “polarization”, which encompasses language that is associated more with one side than the other but is otherwise agnostic about its valence [18]. Similarly, speech may be divisive without being “toxic,” which contains hateful, abusive, or offensive content [19]. Furthermore, personal attacks can be categorized as a form of toxic language, which prior work has examined primarily in the context of online, written communication [20].
In this study, the divisive word lexicon is used to extract specific examples of divisive language used by presidential candidates. The methodology for creating the lexicon is described in §A.5. The final lexicon consisted of 178 words that four researchers reviewed to be qualitatively “divisive” in the context of political speech.
@@
To the best of our knowledge, we build the first divisive speech lexicon. A strength of this lexicon-based analysis is its easy interpretability and applicability. Divisive words are readily identified and may be used by candidates from any political party in a wide variety of settings. However, the measure is inherently limited by the subjectivity in the creation of the lexicon resource and the lack of contextual consideration, similar to existing lexical approaches.
References to political opponents
@@
Once speech referring to opponents is distinguished, we employ the Fightin’ Words method [21] to identify words more strongly associated with opponent mentions; we extend this comparison across multiple candidates by calculating the overlap of each entity’s corresponding word sets with those of all candidates. Intuitively, a greater overlap indicates that other candidates use similar rhetoric in opponent mentions, while a smaller overlap indicates that other candidates do not use similar rhetoric in opponent mentions. This metric thus provides a novel measure for quantifying a candidate’s distinctiveness with respect to their portrayal of opponents by combining lexical and graph analysis. The results are readily interpretable and clarify the distinctive qualities of language used to characterize one’s opponent. The metric is similarly constrained by the pool of candidates and the finite selection of descriptors.
@@ 3 Results
Figure 2: Average sentence uniqueness for each speaker, across all data types. Higher bars indicate greater uniqueness, i.e., that speaker’s speech is less likely to be uttered by other candidates. Trump is the most distinctive speaker among these candidates for debates and SOTU speeches. Error bars correspond to the 95% confidence intervals.
Figure 3: Trump’s average sentence uniqueness scores compared to the others aggregated by party. The number of sentences is denoted on each bar (with error bars representing 95%-confidence intervals). Trump is significantly distinct from other Republicans. The uniqueness scores of the remaining Republicans are closer to those of Democrats than with Trump. For SOTU and campaign speeches, Republicans (without Trump) are less unique than Democrats on average.
@@ 3.1 Donald Trump is unique among all presidential candidates in all types of speech
We start by presenting results based on the LLM-based uniqueness metric. We find that Trump is the most distinctive speaker in debates and SOTU speeches (see Figs. 2 and 3). Additionally, among campaign speeches analyzed from 2008 onwards, Trump speaks in ways that stand apart from all other candidates. Some less distinctive speakers include Bill Clinton in debates and George Bush in SOTU; Jimmy Carter is one of the least distinctive speakers for both debates and SOTU. By comparing candidate speech patterns aggregated by party, Fig. 3 shows that Trump is more distinctive than his fellow Republicans for all types of speech. Indeed, the observed difference between Democratic and Republican candidates is minor compared to the gap observed between Trump and everyone else.
While Trump’s speech is characterized by shorter sentences (Fig. 1), we confirm that Trump’s uniqueness is consistent across sentences of all lengths in Fig. 4. We also note that Biden has similarly short sentences on average as Trump, but his uniquness scores are close in magnitude to those of the other candidates, particularly in debates and SOTU.
@@
Further extensions and robustness checks
@@ 3.2 Trump speaks most divisively
Divisive word lexicon
@@
DEBATES
SOTU
CAMPAIGN
@@
Table 2: Top 10 used divisive words by Trump in each type of speech. Raw counts of usage are denoted in parentheses next to each word.
One common trend shown by the frequency plots for each dataset is Donald Trump’s relatively high usage of words from the divisiveness lexicon compared to other speakers. In all three datasets, Donald Trump’s speeches rank highest in divisive word usage compared to the other candidates analyzed. Examples of Trump’s most frequently used words from this lexicon usage are shown in Table 2. He is most verbosely divisive in debates and campaign speeches, uttering terms like “crazy”, “corrupt”, and “stupid” with high frequency. The contexts of debates and campaigns are more combative than SOTU addresses, which is reflected in the higher levels of divisive language in those two mediums. Additional example sentences from various speakers can be found in §C.
DEBATES
@@
Table 3: Example sentences from debates and campaigns, with preceding context, spoken by Trump that use words from the divisive lexicon.

### v2

# Quantifying the Uniqueness and Divisiveness of Presidential Discourse

@@ Abstract
Do American presidents speak discernibly different from each other? If so, in what ways? And are these differences confined to any single medium of communication? To investigate these questions, this paper introduces a novel metric of uniqueness based on large language models, develops a new lexicon for divisive speech, and presents a framework for assessing the distinctive ways in which presidents speak about their political opponents. Applying these tools to a variety of corpora of presidential speeches, we find considerable evidence that Donald Trump’s speech patterns diverge from those of all major party nominees for the presidency in recent history. Trump is significantly more distinctive than his fellow Republicans, whose uniqueness values appear closer to those of the Democrats. Contributing to these differences is Trump’s employment of divisive and antagonistic language, particularly when targeting his political opponents. These differences hold across a variety of measurement strategies, arise on both the campaign trail and in official presidential addresses, and do not appear to be an artifact of secular changes in presidential communications.
@@
Significance Statement
While presidential discourse attracts considerable attention from popular media and scholars alike, efforts to computationally compare such rhetoric have been limited to lexical methods. This paper proposes a novel suite of metrics to advance such analyses and identify new findings. In particular, we leverage large language models to establish an original metric of uniqueness, develop a new lexicon for divisive speech, and introduce a comparative framework for the portrayal of political opponents. We then apply these methods to a rich assembly of campaign speeches, presidential debates, and official States of the Union addresses. Across all these datasets, we find that Donald Trump’s political rhetoric is unique among modern presidents, and is defined, in part, by his use of antagonistic language, particularly when directed at political opponents.
@@ 1 Introduction
The rise of the modern presidency is defined, in no small measure, by the chief executive’s changing relationship to the American public. In what Jeffrey Tulis [1] calls the “rhetorical presidency,” modern presidents are expected to routinely stand before the public in order to explain, persuade, inform, and instruct. How precisely they do so, though, is open to interpretation. While certain norms of communication govern their behavior, presidents have a fair measure of discretion to speak as they choose.
@@
When fulfilling their oratory duties, do modern presidents adhere to a common script? Or do some presidents defy rhetorical conventions and speak in ways that, at least among themselves, are novel and surprising? Leveraging recent advances in large language models (LLMs) [2], we develop a new quantifier of uniqueness by directly measuring the unpredictability of language patterns. In addition, we incorporate more standard lexical techniques to examine a prevalent yet understudied construct, divisiveness [3]. We operationalize divisiveness as language that is intended to impugn and delegitimize the speaker’s target, and we develop a new lexicon for such speech. Furthermore, we introduce a comparative framework for assessing mentions of opponents, which are especially prevalent in presidential debates.
We use our proposed tools to analyze large and diverse corpora of presidential speech. In doing so, we are able to both characterize the overall distinctiveness of presidents’ speech patterns and examine specific qualities that previous research has overlooked. Moreover, we can distinguish general speech patterns from those directed towards one’s political opponents.
In nearly all of our analyses, Donald Trump appears as a clear outlier. On the campaign trail, in presidential debates, and in official presidential addresses, we find, Trump’s speech patterns routinely differ from those of all recent presidents — lending credence to Kurt Anderson’s observation that, “The version of English [Trump] speaks amounts to its own patois, with a special vocabulary and syntax and psychological substrate” [4].
Like previous scholars [5, 6, 7], we find that Trump tends to communicate in shorter, more simplistic sentences. But whereas previous quantitative research on presidential rhetoric relied exclusively on lexicons (e.g., [7]) and sentiment classification (e.g., [8]), which necessarily disregard contextual information included in the body of speeches, our methods are able to show that Trump speaks in ways that are holistically different from all modern presidents. These differences are pervasive and large — so much so, in fact, that the observed differences between Trump and his fellow Republicans exceed those between Republicans and Democratic presidents. Contributing to these differences, we show, is Trump’s tendency to speak in ways that are especially divisive, particularly when focusing on his political opponents. [^1] This finding is consistent with previous research showing that Trump uses language that is explicitly intended to evoke an emotional response from his revision audience [9], and that Trump’s speech patterns reveal his autocratic ambitions [10], anti-democratic views [11], a willingness to condone political violence [12], a commitment to populism [3], and low levels of analytic thinking [7]. None of these scholars, however, have developed a lexicon of divisiveness, as we do in this paper. These findings, moreover, are robust to a variety of measurement strategies, arise across rich and diverse corpora of texts, and do not appear to be an artifact of secular time trends.
@@ 2.1 Data
Our research investigates three genres of political speech: presidential debates in general elections since 1960, State of the Union (SOTU) speeches since 1961, and a sample of campaign speeches assembled by the American Presidency Project [13]. Debates and SOTU speeches are generally standard across presidents and time, so all available documents are included in our main analyses. [^2] Additional data are used for model training. See §A.3 for details. Publicly available campaign documents, however, are imbalanced and not comprehensive over the same time frame, so these corpora are limited to speeches delivered within one month of Election Day in every presidential election since 2008. Table 1 provides summary statistics of the final datasets for which we present results. [^3] Within these data, Donald Trump spoke in five debates (3,610 sentences), four SOTU addresses (1,471 sentences), and 28 campaign speeches (7,488 sentences) between the years of 2016-2020.
Fig. 1 shows the distribution of sentence lengths from each speaker across the datasets. Further details of how we collected these datasets can be found in §A.1 From the outset, however, we note that Trump tends to speak in markedly shorter sentences than do other presidents. Whereas Trump’s sentences range from 10.4 to 14.5 words in the three data sources, the overall averages range from 17.6 to 24.4 words. And when comparing presidents within each data source, Trump registered the single lowest number of average words per sentence among all presidents within debates and campaigns, and the second lowest number within SOTU addresses.
Figure 1: Distribution of sentence lengths among all speakers, across debates, SOTU, and campaign data (error bars represent 95% CI). Donald Trump tends to use the shortest sentences on average.
@@ 2.2 Quantifying the uniqueness of political speech
We quantify the uniqueness of political speeches using three complementary approaches: (i) a novel metric based on LLMs; (ii) a new resource of lexicons for divisive speech; and (iii) a comparative framework for the portrayal of political opponents involving lexical features. Together, these three approaches allow for a robust and multifaceted comparison of speech patterns that involve global assessments beyond lexical comparisons and specific evaluations of divisiveness. We briefly introduce the intuition behind each approach below.
LLM-based uniqueness.
@@
Large language models, like the GPT family of models, have received widespread attention for their abilities to statistically characterize the complex structures of natural language text. LLMs can measure the predictability of text by calculating the likelihood of the next word or sequence of words in a given context, to produce a measure known as “perplexity,” which is typically used to evaluate the quality of LLMs [14, 15]. To control for the length of a text, standard perplexity measures can be supplemented with measures of “bits-per-character” (BPC) [16].
@@
Contributing to this literature, we propose a metric of “uniqueness” based on the ability of LLMs to estimate the probabilities of word sequences, and we then use these estimates to compare political speech from various presidents and presidential candidates, whether delivered from the White House or on the campaign trail. Specifically, from a pool of presidential candidates, we determine how likely the speech of one speaker is to be produced by the others. A positive value for one candidate’s sentence uniqueness, therefore, suggests that other candidates are unlikely to say it. The larger this value, the less likely other candidates are to do so.
The advantage of this metric is its consideration of the preceding context of given speech. Rather than examine words or phrases in isolation, this metric considers the order in which they appear, and thereby provides a much more nuanced characterization of speech patterns. The precise meaning of the scores, however, may be difficult to interpret because they do not reveal the exact features that make a speech distinctive, and because they only allow for comparisons within a specified pool of speakers. Technical details on the construction of this metric are included in the §A.3 and §A.4.
Divisive speech lexicon.
@@
To analyze the actual content of language used by presidents and presidential candidates, a “divisiveness” lexicon is created and applied to each dataset. We define language as “divisive” if it intends to impugn and delegitimize the speaker’s target, e.g., by attacking their intelligence, integrity, or intentions. Examples of divisive accusations include “racist”, “dishonest”, “corrupt”, or “ridiculous”. Such labels are expressly designed to put the target on defense and accentuate differences and distance between parties.
Our definition of divisive is distinct from other commonly analyzed constructs such as political “polarization”, which encompasses language that is associated more with one side than the other but is otherwise agnostic about its valence [17]. Meanwhile, speech may be divisive without being “toxic,” which contains hateful, abusive, or offensive content [18]. Personal attacks can also be categorized as a form of toxic language, which prior work has examined primarily in the context of online, written communication [19]. Nor is our measure of divisiveness the simple antonym of traditional notions of politeness [20], which usually involve honoring social conventions, showing gratitude, paying compliments, avoiding complaints and curses, and respecting the listener’s autonomy with the use of softening statements and hedges. More than just impolite or insulting, divisive speech, as we conceive it, is explicitly intended to serve the political purposes of delegitimization, marginalization, and distancing between speaker and target.
@@
To the best of our knowledge, ours is the first divisive speech lexicon. As we explain in further detail in §A.5, this lexicon consists of 178 words that four researchers independently reviewed to be qualitatively “divisive” in political speech. A strength of this lexicon-based analysis is its easy interpretability and applicability. Divisive words may be used by candidates from any political party in a wide variety of settings, are readily identified, and, as we subsequently show, are broadly agreed upon by coders. As with all lexical approaches, however, the measure is inherently limited by the subjective nature of lexical evaluations and the lack of contextual consideration. Our resource, like all lexicons, by itself does not account for surrounding context like negation or valence. Strategies like pairing our lexicon with part-of-speech taggers may be leveraged for downstream tasks.
References to political opponents.
@@
Once speech referring to opponents is distinguished, we employ the Fightin’ Words (FW) method [21] to identify words more strongly associated with opponent mentions; we extend this comparison across multiple candidates by calculating the overlap of each entity’s corresponding word sets with those of all presidents and presidential candidates. Intuitively, a greater overlap indicates that other candidates use similar rhetoric in opponent mentions, while a smaller overlap indicates that other candidates do not use similar rhetoric in opponent mentions. This metric thus provides a novel measure for quantifying a candidate’s distinctiveness with respect to their portrayal of opponents by combining lexical and graph analysis. The results are readily interpretable and clarify the distinctive qualities of language used to characterize one’s opponent. The metric is similarly constrained by the pool of candidates and the finite selection of descriptors.
@@ 3.1 Donald Trump is unique among all presidential candidates in all types of speech
We start by presenting results based on the LLM-based uniqueness metric. We find that Trump is the most distinctive speaker in debates and SOTU speeches (see Figs. 2 and 3). Additionally, among campaign speeches analyzed from 2008 onwards, Trump speaks in ways that stand apart from all other candidates. By comparing candidate speech patterns aggregated by party, Fig. 3 shows that Trump is more distinctive than his fellow Republicans for all types of speech. Indeed, the observed difference between Democratic and Republican candidates is minor compared to the gap observed between Trump and everyone else.
Figure 2: Average sentence uniqueness for each speaker, across all data types. Higher bars indicate greater uniqueness, i.e., that speaker’s speech is less likely to be uttered by other candidates. Trump is the most distinctive speaker among these candidates for debates and SOTU speeches. Error bars correspond to the 95% CI.
Figure 3: Trump’s average sentence uniqueness scores compared to the others aggregated by party. The number of sentences is denoted on each bar (with error bars representing 95% CI). Trump is significantly distinct from other Republicans. The uniqueness scores of the remaining Republicans are closer to those of Democrats than with Trump. For SOTU and campaign speeches, Republicans (without Trump) are less unique than Democrats on average.
While Trump’s speech is characterized by shorter sentences (Fig. 1), we confirm that Trump’s uniqueness is consistent across sentences of all lengths in Fig. 4. We also note that Biden has similarly short sentences on average as Trump, but his uniqueness scores are close in magnitude to those of the other candidates, particularly in debates and SOTU addresses.
@@
Further extensions and robustness checks.
@@
Score validation.
The results presented in the main paper are based on the OpenAI GPT-2 model, which was the state-of-the-art open-source causal language model at the start of this project [22]. We are interested in causal language modeling, i.e., the predictability of text in a forward sequential manner. Causal language models are more appropriate for our task of scoring uniqueness, as opposed to masked language models (e.g., BERT, RoBERTa) which are designed to predict missing tokens within a sequence based on bidirectional context (and thus excel at text classification). We validate our results with more recent and powerful LLMs like Gemma 2B [23] and Phi1-5b [24] (see §B.6). The results from these recent LLMs reaffirm our original findings; that is, Trump is consistently identified as the most unique speaker among modern presidential candidates in all types of speech.
@@ 3.2 Trump speaks most divisively
Divisive word lexicon.
@@
Table 2: Top 10 divisive words used by Trump in each type of speech. Raw counts of usage are denoted in parentheses next to each word.
DEBATES SOTU CAMPAIGN
@@
One common trend shown by the frequency plots for each dataset is Donald Trump’s relatively high usage of words from the divisiveness lexicon compared to other speakers. In all three datasets, Donald Trump’s speeches rank highest in divisive word usage compared to the other candidates analyzed. Examples of Trump’s most frequently used words from this lexicon usage are shown in Table 2. He is most verbosely divisive in debates and campaign speeches, uttering terms like “crazy”, “corrupt”, and “stupid” with high frequency. The contexts of debates and campaigns are more combative than SOTU addresses, which is reflected in the higher levels of divisive language in those two mediums. Table 3 contains a selection of Trump’s divisiveness in context; additional example sentences from various speakers can be found in the §C.
Across all debates and campaign speeches, which are the two corpora with greatest prevalence of divisive language, sentences that use divisive words are more unique (see §E). The Spearman correlation coefficient between uniqueness and divisive word usage is 0.01 in debates and campaigns ($p<0.05$ in campaigns and $p=0.30$ in debates). The weak, positive correlation suggests that divisiveness amounts to only a small part of the overall uniqueness of any speech pattern.
Table 3: Example sentences from debates and campaigns, with preceding context, spoken by Trump that use words from the divisive lexicon. Bolded words are matched with our divisiveness lexicon.
DEBATES: